OpenStudy (anonymous):

(HELP!!!) Simplify the radical expression by rationalizing the denominator. [4√(150)/(√189x)]

OpenStudy (anonymous):

\[\frac{ 4\sqrt{150} }{ \sqrt{189x} } = \frac{ 4\sqrt{150}\sqrt{189x} }{ 189x }\]

OpenStudy (anonymous):

\[4\sqrt{150}\sqrt{189x} = 180\sqrt{14}\sqrt{x}\] all over 189x

OpenStudy (anonymous):

That doesn't make any sense.

OpenStudy (anonymous):

4(sqrt(150) = 20sqrt(6) 20sqrt(6) * sqrt(189x) = 180sqrt(14) sqr(x) The above is not fully simplified

OpenStudy (anonymous):

\[\sqrt{189x} = \sqrt{189} * \sqrt{x}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

4sqrt(150): sqrt(100) = 10 sqrt(121) = 11 sqrt(141) = 12 sqrt(150) = 12.247.... 12.247.. = 5sqrt(6) 4sqrt(150) = 4(5(sqrt(6)) = 20sqrt(6)